# A note on real algebraic maps which are topologically special generic maps

### 21 Jan 2024 v3

@@ Abstract.
In such stories, we have previously constructed real algebraic maps topologically regarded as special generic maps. This paper is a kind of additional short remark on such maps.
@@
For example, [21, 35, 44] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are simplest symmetric spaces. We are interested in another generalizations of projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [11], followed by [15] by the author himself. We present results of a similar type which are also new.
@@
Here, a real algebraic manifold means some union of connected components of the intersection of finitely many sets represented as the zero sets of some real polynomial maps. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere give simplest examples. Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections. Note that we may say Nash manifolds and maps for example for such maps. However, we do not use such terminologies. [41] show related fundamental theory, arguments and some advanced theory.
@@
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result as Main Theorem 1. This is proven in the next section with additional results Main Theorems 2 and 3. The third section is on meanings of our main result and related topics.
@@
For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
@@
    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$.
@@
    There exists a pair $(\Phi;M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter we can abuse the notation in Proposition 1 and do.
@@
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ are isomorphic and naturally identified.
@@
We add the following as additional main results.
@@
We can have a pair of maps in Main Theorem 1 as follows.
  1. (1)
    $f$ is product-organized.
  2. (2)
    $M_{0}\subset{\mathbb{R}}^{m+1}$ is also represented as the zero set of some real polynomial.
Main Theorem 3.
In Main Theorem 2, we can have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
  1. (1)
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$, the boundary of whose fiber is diffeomorphic to $S^{m-n}$ and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$: in fact a smooth manifold which is homeomorphic to $D^{k}$ and whose fiber is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined before except for $m-n+1=k\neq 4$.
  2. (2)
    Under the presented structure group, the bundle of (1) here is in fact trivial.
  3. (3)
    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ diffeomorphic to a manifold obtained in the following steps.
    1. (a)
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is diffeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
    2. (b)
    3. (c)
      Eliminate the corner.
    In fact, except the case where ${\Delta}^{k}$ is not diffeomorphic to $D^{k}$, $M$ and $M_{0}$ are diffeomorphic.
Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$.
A proof of Main Theorems.
First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
@@
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We explain about this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then for some variable $x_{j_{0}}$, we have the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$. The first value is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class. The second value is not zero by the property of the zero sets of the polynomials.
@@
is a smooth, closed and connected manifold and non-singular real algebraic manifold of dimension $m$. Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
STEP 3 Constructing a special generic map $f$.
@@
For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
We construct a homeomorphism from $M_{0}$ onto another smooth manifold $M$. First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$ where $0\in{\mathbb{R}}^{m-n+1}$ here is the origin of course.
We can consider a smooth function $g:(0,\infty]\rightarrow(0,1]$ enjoying the following.
@@
  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$
@@
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the canonical way.
@@
We have a homeomorphism onto another manifold $M$. This is smooth and homeomorphic to $M_{0}$. We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
The resulting map is also special generic by the construction. It is also product-organized. By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2.
This completes the proof of Main Theorems 1 and 2.
We show Main Theorem 3. We can prove this by applying (revised versions of) known arguments previously presented in related studies and our present study.
We define $W_{0}$ as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$
and this can be also defined as
$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$.
We use the function $g$ before. We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule where we consider some elements as vectors.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
@@
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$.

### 15 Feb 2024 v4

# Some remarks on real algebraic maps which are topologically special generic maps

@@ Abstract.
In such stories, we have previously constructed some real algebraic maps topologically regarded as special generic maps. This paper constructs such maps in general situations and is mainly for additional remarks on such maps.
@@
For example, [24, 39, 50] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are simplest symmetric spaces. We are interested in another generalizations of projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [14], followed by [17] by the author himself. We present new related results: related new examples.
@@
Here, a real algebraic manifold means a union of connected components of the zero sets of a real polynomial map. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere give simplest examples. Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections. Note that we may say Nash manifolds and maps for example for such maps. However, we do not use such terminologies. [45] show related fundamental theory, arguments and some advanced theory.
@@
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result as Main Theorem 1. This is proven in the next section with additional results Main Theorems 2 and 3. Main Theorem 3 is a more generalized result. The third section is on meanings of our main result and related topics.
@@
Let $m>n\geq 1$ be integers. For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
@@
    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ which is also the zero set of a real polynomial.
@@
    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$.
This paper is regarded as an improved version of [18] and based on our new discovery and idea.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter we can abuse the notation in Proposition 1 and do. We can also abuse this in the case $m=n$ and we do.
@@
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified.
@@
Note that in Proposition 1, we cannot set ”$m\geq n$”. This is closely related to Eliashberg’s celebrated theory [6]. See also [7].
@@
We add the following as another additional main result.
@@
In Main Theorem 1, $f$ and $f_{0}$ can be obtained as product-organized special generic maps.
A proof of Main Theorems 1 and 2.
STEPs 1–3 are main ingredients.
@@
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We explain about this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then for some variable $x_{j_{0}}$, we have the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$. The value of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ for $x_{j_{0}}$ is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class. The product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$ is not zero by the property of the zero sets of the polynomials.
@@
is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$. Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
STEP 3 Is $f_{0}$ special generic?
We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $\partial{\bar{N}}_{0}\times[0,1]\subset N(\partial{\bar{N}}_{0})$. More precisely, we can have one and consider identification in the following way thanks to arguments related to ”Definition 14” [20].
  • $\partial{\bar{N}}_{0}\times\{0\}$ is identified with $\partial N(\partial{\bar{N}}_{0})\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
  • $\partial{\bar{N}}_{0}\times\{t_{0}\}\subset\partial{\bar{N}}_{0}\times(0,1]$ is identified with the zero set of the polynomial $f_{j}-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function: some $y_{j}$ is regarded as the variable of the target of the function.
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a diffeomorphism from ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ onto ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ and by the restriction another diffeomorphism $({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto ${\rm Int}\ {\bar{N}}_{0}\times S^{m-n}\subset{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
These arguments imply that $f_{0}$ is a product-organized special generic map.
We argue STEPs 1–3 again in a generalized way in the proof of Main Theorem 3. By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
This completes the proof.
Main Theorem 3.
Suppose a situation as in Main Theorem 1 is given.
Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following. We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $t$.
  • $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
  • $F_{0,S^{m-n}}(0)=0$. $F_{0,S^{m-n}}(t)>0$ for $t>0$ and $F_{0,S^{m-n}}(t)<0$ for $t<0$. The value of the derivative of $F_{0,S^{m-n}}$ at 00 is not zero.
  • ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$. $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)\geq 0$ for $y\in{\mathbb{R}}^{m-n+1}$.
  • ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the empty set for $t<0$.
  • For $a_{1},a_{2}\in\mathbb{R}$ such that $a_{1}<a_{2}$, define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$. We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)$ mapping ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively. ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
  • For any real number $\epsilon>0$, we have a positive real number $\delta>0$ and for $0<t<\delta$ and each point $p$ of ${F_{S^{m-n},t}}^{-1}(0)$, $0<||p||<\epsilon$.
Then we can have a statement of Main Theorem 1 by defining
$M_{0}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
for a sufficiently large $T>0$.
As a more specific case, we can also have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
  1. (1)
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$. This is also a bundle which may not be smooth and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$.
  2. (2)
    Under the presented structure group, the bundle of (1) here which may not be a smooth bundle is in fact trivial.
  3. (3)
    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps.
    1. (a)
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is homeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4,5$. Note also that for $k\neq 4$ this is diffeomorphic to $D^{k}$ if its boundary is diffeomorphic to $S^{k-1}$.
    2. (b)
    3. (c)
      Eliminate the corner.
We review several known facts on topologies and differentiable structures of spheres and disks shortly. See [34] and see also [12, 46, 47] for example. A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example. It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$. If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$. A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered. A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$. A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$. Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
A proof of Main Theorem 3.
We can argue as STEP 1 of the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than $a$. We choose another positive number $T_{0}<a$.
We present an argument similar to one of STEP 2 of the proof of Main Theorems 1 and 2. Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
This mainly comes from the following assumption: ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 of the proof of Main Theorems 1 and 2. The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not zero from the assumption, and the value in the case of Main Theorems 1 and 2. The resulting value is not zero. We can argue similarly to prove that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
@@
We present an argument similar to one of STEP 3 of the proof of Main Theorems 1 and 2. For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
We construct a homeomorphism from the manifold $M_{0}$ onto another smooth manifold $M$.
First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$.
Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a,1}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component and ${\rm Pr}_{{\Sigma}^{m-n},a,2}:{\Sigma}^{m-n}\times(0,a]\rightarrow(0,a]$ the projection to the second component.
We can consider a smooth function $g:(0,\infty)\rightarrow(0,1]$ enjoying the following.
@@
  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$.
@@
We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
$(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),T_{0}))$
according to the following rule.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
We have a homeomorphism onto a smooth manifold $M$. We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$. The resulting map is also special generic by the construction. It is also product-organized.
We define $W_{0}$ as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
and this can be also defined as
$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$.
We use the function $g$ before. We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. We thus have a homeomorphism from $W_{0}$. The image of the homeomorphism is the union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
@@
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
We can define another non-singular real algebraic manifold
$M_{0,t}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t)=0\}$
for $0<t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. This is diffeomorphic to $M_{0}$. This new set has properties similar to important ones of $M_{0}$.
We go back to the trivial bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle. This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, this smooth bundle which may not be trivial in general is also trivial. Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
We investigate the bundle in Main Theorem 3 (1) and (2).